An ideal gas at 0 degrees C consists of 1.0*10^23 atoms. 22J of thermal energy are added to the gas.What is the new temperature in degrees C ?
Answer #1
Concepts and reason

The concept of change in internal energy is required to solve the problem.

First, calculate the number of moles by using the number of atoms and Avogadro number. Finally, use the expression of change in internal energy and determine the new temperature of the gas.

Fundamentals

The change in internal energy is given by the relation,

ΔU=32nRΔT\Delta U = \frac{3}{2}nR\Delta T

Here, n is the number of moles, R is the gas constant, and ΔT\Delta T is the change in temperature.

The change in temperature is given as,

ΔT=T2T1\Delta T = {T_2} - {T_1}

Here, T1{T_1} is the initial temperature and T2{T_2} is the final temperature.

The number of moles of a gas is determined by using the relation,

n=NNAn = \frac{N}{{{N_{\rm{A}}}}}

Here, N is the number of atoms and NA{N_{\rm{A}}} is the Avogadro number.

Calculate the number of moles.

The number of moles of the gas is calculated by using the relation,

n=NNAn = \frac{N}{{{N_{\rm{A}}}}}

Here, N is the number of atoms and NA{N_{\rm{A}}} is the Avogadro number.

Substitute 1.0×1023atoms1.0 \times {10^{23}}{\rm{ atoms}} and 6.023×1023moles/atoms6.023 \times {10^{23}}{\rm{ moles/atoms}} in the above equation and solve for n.

n=(1.0×1023atoms)(6.023×1023atoms/moles)=0.166mol\begin{array}{c}\\n = \frac{{\left( {1.0 \times {{10}^{23}}{\rm{ atoms}}} \right)}}{{\left( {6.023 \times {{10}^{23}}{\rm{ atoms/moles}}} \right)}}\\\\ = 0.166{\rm{ mol}}\\\end{array}

The change in internal energy of the gas is given as,

ΔU=32nRΔT\Delta U = \frac{3}{2}nR\Delta T

Here, n is the number of moles of the gas and ΔT\Delta T is the change in temperature of the gas.

Substitute T2T1{T_2} - {T_1} for ΔT\Delta T in the above equation ΔU=32nRΔT\Delta U = \frac{3}{2}nR\Delta T and solve for T2{T_2} .

ΔU=32nR(T2T1)T2T1=2ΔU3nRT2=2ΔU3nR+T1\begin{array}{c}\\\Delta U = \frac{3}{2}nR\left( {{T_2} - {T_1}} \right)\\\\{T_2} - {T_1} = \frac{{2\Delta U}}{{3nR}}\\\\{T_2} = \frac{{2\Delta U}}{{3nR}} + {T_1}\\\end{array}

Here, T1{T_1} is the initial temperature of the gas and T2{T_2} is the new temperature of the gas.

Substitute 22 J for ΔU\Delta U , 0oC0{{\rm{ }}^{\rm{o}}}{\rm{C}} for T1{T_1} , 0.166 moles for n, and 8.314J/moloC8.314{\rm{ J/mol}}{ \cdot ^{\rm{o}}}{\rm{C}} for R in the equation T2=2ΔU3nR+T1{T_2} = \frac{{2\Delta U}}{{3nR}} + {T_1} .

T2=2(22J)3(0.166mol)(8.314J/moloC)+0oC=10.6oC\begin{array}{c}\\{T_2} = \frac{{2\left( {22\;{\rm{J}}} \right)}}{{3\left( {0.166{\rm{ mol}}} \right)\left( {8.314{\rm{ J/mol}}{ \cdot ^{\rm{o}}}{\rm{C}}} \right)}} + 0{{\rm{ }}^{\rm{o}}}{\rm{C}}\\\\{\rm{ = }}10.6{{\rm{ }}^{\rm{o}}}{\rm{C}}\\\end{array}

Ans:

The new temperature of the gas is 10.6oC10.6{{\rm{ }}^{\rm{o}}}{\rm{C}} .